### v1

@@ 6 Formal Analysis of Some Core Methods
Resampling or cost-sensitive loss minimization. A common approach for handling imbalanced data in practice is to assign distinct costs to positive and negative samples. This technique, implemented either explicitly or through resampling, is widely used in empirical studies (Chawla et al., 2002; He and Garcia, 2009; He and Ma, 2013; Huang et al., 2016; Buda et al., 2018; Cui et al., 2019). The associated target loss ${\mathsf{L}}_{c_{+},c_{-}}(h,x,y)$ can be expressed as follows, for any $c_{+}>0$, $c_{-}>0$ and $(h,x,y)\in{\mathscr{H}}_{\mathrm{all}}\times{\mathscr{X}}\times{\mathscr{Y}}$:
@@
$\rho_{+}=\frac{2m^{\frac{1}{4}}_{-}}{m^{\frac{1}{4}}_{+}+m^{\frac{1}{4}}_{-}}\rho_{\rm{geom}}$ $\rho_{-}=\frac{2m^{\frac{1}{4}}_{+}}{m^{\frac{1}{4}}_{+}+m^{\frac{1}{4}}_{-}}\rho_{\rm{geom}}.$
Thus, disregarding the suboptimal exponent of $\frac{1}{4}$ compared to $\frac{1}{3}$, which results from a less precise technical analysis, the margin values recommended in their work directly contradict those suggested by our analysis, see Eqn. (2). Specifically, their analysis advocates for a smaller positive margin when $m_{+}>m_{-}$, whereas our theoretical analysis prescribes the opposite. This discrepancy stems from the analysis in (Cao et al., 2019), which focuses on a balanced loss (a uniform average over positively and negatively labeled points), which deviates fundamentally from the standard zero-one loss we consider. Figure 1 illustrates these contrasting solutions in a specific case of separable data. On the standard zero-one loss, our approach obtains a lower error.
@@ 7 Experiments
In this section, we present experimental results for our immax algorithm, comparing it to baseline methods in minimizing the standard zero-one misclassification loss on CIFAR-10, CIFAR-100 (Krizhevsky, 2009) and Tiny ImageNet (Le and Yang, 2015) datasets. .
Starting with multi-class classification, we strictly followed the experimental setup of Cao et al. (2019), adopting the same training procedure and neural network architectures. Specifically, we used ResNet-34 with ReLU activations (He et al., 2016), where ResNet-$n$ denotes a residual network with $n$ convolutional layers. For CIFAR-10 and CIFAR-100, we applied standard data augmentations, including 4-pixel padding followed by $32\times 32$ random crops and random horizontal flips. For Tiny ImageNet, we used 8-pixel padding followed by $64\times 64$ random crops. All models were trained using Stochastic Gradient Descent (SGD) with Nesterov momentum (Nesterov, 1983), a batch size of $1,024$, and a weight decay of $1\times 10^{-3}$. Training spanned $200$ epochs, using a cosine decay learning rate schedule (Loshchilov and Hutter, 2016) without restarts, with the initial learning rate set to $0.2$. For all the baselines and the immax algorithm, the hyperparameters were selected through cross-validation.
To create imbalanced versions of the datasets, we reduced the percent of examples per class identically in the training and test sets. Following (Cao et al., 2019), we consider two types of imbalances: long-tailed imbalance (Cui et al., 2019) and step imbalance (Buda et al., 2018). The imbalance ratio, $\rho=\frac{\max_{k=1}^{c}m_{k}}{\min_{k=1}^{c}m_{k}}$, represents the ratio of sample sizes between the most frequent and least frequent classes. In the long-tailed imbalance setting, class sample sizes decrease exponentially across classes. In the step setting, minority classes all have the same sample size, as do the frequent classes, creating a clear distinction between the two groups.
We compare our immax algorithm with widely used baselines, including the cross-entropy (ce) loss, Re-Weighting (rw) method (Xie and Manski, 1989; Morik et al., 1999), Balanced Softmax (bs) loss (Jiawei et al., 2020), Equalization loss (Tan et al., 2020), Logit Adjusted (la) loss (Menon et al., 2021), Class-Balanced (cb) loss (Cui et al., 2019), the focal loss in (Ross and Dollár, 2017) and the ldam loss in (Cao et al., 2019) detailed in Appendix B. We average accuracies on the imbalanced test set over five runs and report the means and standard deviations. Experimental details on cross-validation are provided in Appendix B. Note that immax is not optimized for other objectives, such as the balanced loss, and thus is not expected to outperform state-of-the-art methods tailored to those metrics.
@@
Let us emphasize that our work is based on a novel, principled surrogate loss function designed for imbalanced data. Accordingly, we compare our new loss function directly against existing ones without incorporating additional techniques. However, all these loss functions, including ours, can be combined with existing data modification methods such as oversampling (Chawla et al., 2002) and undersampling (Wallace et al., 2011; Kubat and Matwin, 1997), as well as optimization strategies like the deferred re-balancing schedule proposed in (Cao et al., 2019), to further enhance performance. For a fair comparison of loss functions, we deliberately excluded these techniques from our experiments.
@@ Appendix A Related Work
The class imbalance problem, defined by a significant disparity in the number of instances across classes within a dataset, is a common challenge in machine learning applications (Lewis and Gale, 1994; Fawcett and Provost, 1996; Kubat and Matwin, 1997; Kang et al., 2021; Menon et al., 2021; Liu et al., 2019; Cui et al., 2019). This issue is prevalent in many real-world binary classification scenarios, and arguably even more so in multi-class problems with numerous classes. In such cases, a few majority classes often dominate the dataset, leading to a “long-tailed” distribution. Classifiers trained on these imbalanced datasets often struggle, performing similarly to a naive baseline that simply predicts the majority class.
The problem has been widely studied in the literature (Cardie and Nowe, 1997; Kubat and Matwin, 1997; Chawla et al., 2002; He and Garcia, 2009; Wallace et al., 2011). It includes numerous methods including standard Softmax, class-sensitive learning, Weighted Softmax, weighted 0/1 loss (Gabidolla et al., 2024), size-invariant metrics for Imbalanced Multi-object Salient Object Detection studied by Li et al. (2024a) as well as Focal loss (Lin et al., 2017), LDAM (Cao et al., 2019), ESQL (Tan et al., 2020), Balanced Softmax (Jiawei et al., 2020), LADE (Hong et al., 2021)), logit adjustment (UNO-IC (Tian et al., 2020), LSC (Wei et al., 2024)), transfer learning (SSP (Yang and Xu, 2020)), data augmentation (RSG (Wang et al., 2021a), BSGAL (Zhu et al., 2024), ELTA (Liu et al., 2024), OT (Gao et al., 2023)), representation learning (OLTR (Liu et al., 2019), PaCo (Cui et al., 2021), DisA (Gao et al., 2024), RichSem (Meng et al., 2023), RBL (Meng et al., 2023), WCDAS (Han, 2023)), classifier design (De-confound (Tang et al., 2020), (Yang et al., 2022; Kasarla et al., 2022), LIFT (Shi et al., 2024), SimPro (Du et al., 2024)), decoupled training (Decouple-IB-CRT (Kang et al., 2020), CB-CRT (Kang et al., 2020), SR-CRT (Kang et al., 2020), PB-CRT (Kang et al., 2020), MiSLAS (Zhong et al., 2021)), ensemble learning (BBN (Zhou et al., 2020), LFME (Xiang et al., 2020), RIDE (Wang et al., 2021b), ResLT (Cui et al., 2022), SADE (Zhang et al., 2022a), DirMixE (Yang et al., 2024)). An interesting recent study characterizes the asymptotic performances of linear classifiers trained on imbalanced datasets for different metrics (Loffredo et al., 2024).
@@
Data modification methods. These include methods such as oversampling the minority class (Chawla et al., 2002), undersampling the majority class (Wallace et al., 2011; Kubat and Matwin, 1997), or generating synthetic samples (e.g., SMOTE (Chawla et al., 2002; Qiao and Liu, 2008; Han et al., 2005)), aim to rebalance the dataset before training (Chawla et al., 2002; Estabrooks et al., 2004; Liu et al., 2008; Zhang and Pfister, 2021; Shi et al., 2023).
@@
Cost-sensitive techniques. These techniques, including cost-sensitive learning and the incorporation of class weights assign different penalization costs to losses on different classes. They include cost-sensitive SVM (Iranmehr et al., 2019; Masnadi-Shirazi and Vasconcelos, 2010) and other cost-senstive methods (Elkan, 2001; Zhou and Liu, 2005; Zhao et al., 2018; Zhang et al., 2018, 2019; Sun et al., 2007; Fan et al., 2017; Jamal et al., 2020; Zhang et al., 2022b; Wang et al., 2022; Xiao et al., 2023; Suh and Seo, 2023). The weights are often determined by the relative number of samples in each class or a notion of effective sample size Cui et al. (2019).
@@
These two method categories are very related and can actually be shown to be equivalent in the limit. Cost-sensitive methods can be viewed as more efficient, flexible and principled techniques for implementing data sampling methods. However, these methods often risk overfitting the minority class or discarding valuable information from the majority class. Both methods inherently bias the input training data distribution and suffer from Bayes inconsistency (in Section, we prove that cost-sensitive methods do not admit Bayes consistency). While they have been both reported to be effective in various instances, this varies and depends on the problem, the distribution, the choice of predictors, and the performance metric adopted and they have been reported not to be effective in all cases (Van Hulse et al., 2007). Additionally, cost-sensitive methods often resort to careful tuning of hyperparameters. Hybrid approaches attempt to combine the strengths of data modification and cost-sensitive methods but often inherit their respective limitations.
@@
Logistic loss modifications. A family of more recent methods rely on logistic loss modifications. They consist of modifying the logistic loss by augmenting each logit (or predicted score) with an additive hyperparameter. They can be equivalently described as a cost-sensitive modification of the exponential terms appearing in the definition of the logistic loss. They include the Balanced Softmax loss Jiawei et al. (2020), the Equalization loss Tan et al. (2020), and the ldam loss Cao et al. (2019). Other similar additive change methods use quadratically many hyperparameters with a distinct additive parameter for each pair of logits. They include the logit adjustment methods of Menon et al. (2021) and Khan et al. (2019). Menon et al. (2021) argue that their specific choice of the hyperparameter values is Bayes-consistent. A multiplicative modification of the logits, with one hyperparameter per class label is advocated by Ye et al. (2020). This can be equivalently viewed as normalizing scoring functions (or feature vectors in the linear case) beforehand, which is a standard method used in many learning applications, irrespective of the presence of imbalanced classes. The Vector-Scaling loss of Kini et al. (2021) combines the additive modification of the logits with this multiplicative change. These authors further present an analysis of this method in the case of linear predictors, underscoring the specific benefits of the multiplicative changes. As already pointed out, the multiplicative changes coincide with prior rescaling or renormalization of the feature vectors, however.
@@
Other methods. Additional approaches for tackling imbalanced datasets (see Zhang et al. (2023)) include post-hoc correction of decision thresholds (Fawcett and Provost, 1996; Collell et al., 2016) or weights (Kang et al., 2020; Kim and Kim, 2019)], as well as information and data augmentation via transfer learning, or distillation (Li et al., 2024b).
@@
First, most existing solutions are heuristic-driven and lack a solid theoretical foundation, making their performance difficult to predict across varying contexts. In fact, we are not aware of any analysis of the generalization guarantees for these methods, with the exception of that of Cao et al. (2019). However, as further discussed in Section 6, the analysis presented by these authors is limited to the balanced loss, that is the uniform average of the misclassification on each class. More specifically, their analysis is limited to binary classification and only for the separable case. The balanced loss function differs from the target misclassification loss. It has been argued, and that is important, that the balanced loss admits beneficial fairness properties when class labels correlate with demographic attributes as it treats all class errors equally. The balanced loss is also the metric considered in the analysis of several of the logistic loss modifications papers (Cao et al., 2019; Menon et al., 2021; Ye et al., 2020; Kini et al., 2021). However, class labels do not alway relate to demographic attributes. Furthermore, many other criteria are considered for fairness purposes and in many machine learning applications, the misclassification remains the key target loss function to minimize. We will show that, even in the special case of the analysis of Cao et al. (2019), the solution they propose is the opposite of the one corresponding to our theoretical analysis for the standard misclassification loss. We further show that their solution is empirically outperformed by ours.

### v2

@@ 6 Formal Analysis of Some Core Methods
Resampling or cost-sensitive loss minimization. A common approach for handling imbalanced data in practice is to assign distinct costs to positive and negative samples. This technique, implemented either explicitly or through resampling, is widely used in empirical studies (Chawla et al., 2002; He & Garcia, 2009; He & Ma, 2013; Huang et al., 2016; Buda et al., 2018; Cui et al., 2019). The associated target loss ${\mathsf{L}}_{c_{+},c_{-}}(h,x,y)$ can be expressed as follows, for any $c_{+}>0$, $c_{-}>0$ and $(h,x,y)\in{\mathscr{H}}_{\mathrm{all}}\times{\mathscr{X}}\times{\mathscr{Y}}$:
@@
$\rho_{+}=\frac{2m^{\frac{1}{4}}_{-}}{m^{\frac{1}{4}}_{+}+m^{\frac{1}{4}}_{-}}\rho_{\rm{geom}},$ $\rho_{-}=\frac{2m^{\frac{1}{4}}_{+}}{m^{\frac{1}{4}}_{+}+m^{\frac{1}{4}}_{-}}\rho_{\rm{geom}}.$
Thus, disregarding the suboptimal exponent of $\frac{1}{4}$ compared to $\frac{1}{3}$, which results from a less precise technical analysis, the margin values recommended in their work directly contradict those suggested by our analysis, see Eqn. (3). Specifically, their analysis advocates for a smaller positive margin when $m_{+}>m_{-}$, whereas our theoretical analysis prescribes the opposite. This discrepancy stems from the analysis in (Cao et al., 2019), which focuses on the balanced loss, which equalizes the impact of each class by weighting errors inversely to class frequency, and deviates fundamentally from the standard zero-one loss we consider. Figure 1 illustrates these contrasting solutions in a specific case of separable data. On the standard zero-one loss, our approach obtains a lower error.
@@ 7 Experiments
In this section, we present experimental results for our immax algorithm, comparing it to baseline methods in minimizing the standard zero-one misclassification loss on CIFAR-10, CIFAR-100 (Krizhevsky, 2009) and Tiny ImageNet (Le & Yang, 2015) datasets. .
Starting with multi-class classification, we strictly followed the experimental setup of Cao et al. (2019), adopting the same training procedure and neural network architectures. Specifically, we used ResNet-34 with ReLU activations (He et al., 2016), where ResNet-$n$ denotes a residual network with $n$ convolutional layers. For CIFAR-10 and CIFAR-100, we applied standard data augmentations, including 4-pixel padding followed by $32\times 32$ random crops and random horizontal flips. For Tiny ImageNet, we used 8-pixel padding followed by $64\times 64$ random crops. All models were trained using Stochastic Gradient Descent (SGD) with Nesterov momentum (Nesterov, 1983), a batch size of $1,024$, and a weight decay of $1\times 10^{-3}$. Training spanned $200$ epochs, using a cosine decay learning rate schedule (Loshchilov & Hutter, 2022) without restarts, with the initial learning rate set to $0.2$. For all the baselines and the immax algorithm, the hyperparameters were selected through cross-validation, see Appendix B for details.
To create imbalanced versions of the datasets, we reduced the percent of examples per class identically in the training and test sets. Following (Cao et al., 2019), we consider two types of imbalances: long-tailed imbalance (Cui et al., 2019) and step imbalance (Buda et al., 2018). The imbalance ratio, $\frac{\max_{k=1}^{c}m_{k}}{\min_{k=1}^{c}m_{k}}$, represents the ratio of sample sizes between the most frequent and least frequent classes. In the long-tailed imbalance setting, class sample sizes decrease exponentially across classes. In the step setting, minority classes all have the same sample size, as do the frequent classes, creating a clear distinction between the two groups.
We compare our immax algorithm with widely used baselines, including the cross-entropy (ce) loss, Re-Weighting (rw) method (Xie & Manski, 1989; Morik et al., 1999), Balanced Softmax (bs) loss (Jiawei et al., 2020), Equalization loss (Tan et al., 2020), Logit Adjusted (la) loss (Menon et al., 2021), Class-Balanced (cb) loss (Cui et al., 2019), the focal loss in (Ross & Dollár, 2017) and the ldam loss in (Cao et al., 2019) also detailed in Appendix B. We average accuracies on the imbalanced test set over five runs and report the means and standard deviations. Note that immax is not optimized for other objectives, such as the balanced loss, and thus is not expected to outperform state-of-the-art methods tailored to those metrics.
@@
Let us emphasize that our work is based on a novel, principled surrogate loss function designed for imbalanced data. Accordingly, we compare our new loss function directly against existing ones without incorporating additional techniques. However, all these loss functions, including ours, can be combined with existing data modification methods such as oversampling (Chawla et al., 2002) and undersampling (Wallace et al., 2011; Kubat & Matwin, 1997), as well as optimization strategies like the deferred re-balancing schedule proposed in (Cao et al., 2019), to further enhance performance. For a fair comparison of loss functions, we deliberately excluded these techniques from our experiments.
8 Discussion
Balanced accuracy. Our work focuses on the standard and unmodified zero-one misclassification loss, which remains the primary objective in many machine learning applications, as discussed in Section 1. Accordingly, we report standard accuracy based on this loss function in Section 7. In contrast, some previous studies (e.g., Cao et al. (2019)) report “balanced accuracy”, which equalizes the impact of each class by weighting errors inversely to class frequency.
Selection of $\rho_{k}$. As noted in Section 5, while $\rho_{k}$ can be freely tuned, the search can be guided by the vector $[\rho_{k}/\rho]_{k}$ near ${\mathsf{r}}$, the optimal values in the separable case, a strategy we used in our experiments. For large multi-class settings, we further reduced the search space by assigning shared $\rho_{k}$ to rare classes and distinct ones to frequent classes, improving practicality with minimal performance impact.
Balanced data. immax extends several classical algorithms, such as SVM (Cortes & Vapnik, 1995) and logistic regression (Verhulst, 1838), to the more general setting of imbalanced data. When the training dataset is balanced, the theoretically optimal values of $\rho_{k}$ coincide across all classes, and immax reduces to standard methods. For example, with the logistic loss, it recovers the standard softmax cross-entropy loss with a suitable regularization parameter, thus yielding the same performance.
Large-scale datasets. The dependency of our solution on sample size and dimensionality is similar to that of standard neural networks trained with cross-entropy loss (that is the logistic loss when softmax is applied to logits). Thus, our approach remains practical for large-scale datasets when using optimizers such as SGD (Bottou, 2010), AdaGrad (Duchi et al., 2011) or Adam (Kingma & Ba, 2015). Our solution does depend on the number of classes, but this dependency is inherent to standard multi-class neural network solutions as well.
Contrastive loss and temperature scaling. The form of our immax loss function has some similarity with supervised contrastive losses (e.g., Khosla et al. (2020)), where a scalar temperature parameter is used in the inner product argument of the exponential. However, in our case, distinct parameters are introduced to allow different confidence margins across classes, serving a different purpose than in contrastive learning. Nevertheless, our margin analysis could provide a useful tool for analyzing contrastive learning. Alternatively, the immax loss can be viewed as a form of class-dependent temperature scaling derived from the logistic loss. In fact, our theoretical framework offers insight into the role of temperature parameters more broadly.
9 Conclusion
@@
Acknowledgements
We thank the anonymous reviewers for their valuable feedback and constructive suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Related Work
The class imbalance problem, defined by a significant disparity in the number of instances across classes within a dataset, is a common challenge in machine learning applications (Lewis & Gale, 1994; Fawcett & Provost, 1996; Kubat & Matwin, 1997; Kang et al., 2021; Menon et al., 2021; Liu et al., 2019; Cui et al., 2019). This issue is prevalent in many real-world binary classification scenarios, and arguably even more so in multi-class problems with numerous classes. In such cases, a few majority classes often dominate the dataset, leading to a “long-tailed” distribution. Classifiers trained on these imbalanced datasets often struggle, performing similarly to a naive baseline that simply predicts the majority class.
The problem has been widely studied in the literature (Cardie & Nowe, 1997; Kubat & Matwin, 1997; Chawla et al., 2002; He & Garcia, 2009; Wallace et al., 2011). It includes numerous methods including standard Softmax, class-sensitive learning, Weighted Softmax, weighted 0/1 loss (Gabidolla et al., 2024), size-invariant metrics for Imbalanced Multi-object Salient Object Detection studied by Li et al. (2024a) as well as Focal loss (Lin et al., 2017), LDAM (Cao et al., 2019), ESQL (Tan et al., 2020), Balanced Softmax (Jiawei et al., 2020), LADE (Hong et al., 2021)), logit adjustment (UNO-IC (Tian et al., 2020), LSC (Wei et al., 2024)), transfer learning (SSP (Yang & Xu, 2020)), data augmentation (RSG (Wang et al., 2021a), BSGAL (Zhu et al., 2024), ELTA (Liu et al., 2024), OT (Gao et al., 2023)), representation learning (OLTR (Liu et al., 2019), PaCo (Cui et al., 2021), DisA (Gao et al., 2024), RichSem (Meng et al., 2023), RBL (Meng et al., 2023), WCDAS (Han, 2023)), classifier design (De-confound (Tang et al., 2020), (Yang et al., 2022; Kasarla et al., 2022), LIFT (Shi et al., 2024), SimPro (Du et al., 2024)), decoupled training (Decouple-IB-CRT (Kang et al., 2020), CB-CRT (Kang et al., 2020), SR-CRT (Kang et al., 2020), PB-CRT (Kang et al., 2020), MiSLAS (Zhong et al., 2021)), ensemble learning (BBN (Zhou et al., 2020), LFME (Xiang et al., 2020), RIDE (Wang et al., 2021b), ResLT (Cui et al., 2022), SADE (Zhang et al., 2022), DirMixE (Yang et al., 2024)). An interesting recent study characterizes the asymptotic performances of linear classifiers trained on imbalanced datasets for different metrics (Loffredo et al., 2024).
@@
Data modification methods. These include methods such as oversampling the minority class (Chawla et al., 2002), undersampling the majority class (Wallace et al., 2011; Kubat & Matwin, 1997), or generating synthetic samples (e.g., SMOTE (Chawla et al., 2002; Qiao & Liu, 2008; Han et al., 2005)), aim to rebalance the dataset before training (Chawla et al., 2002; Estabrooks et al., 2004; Liu et al., 2008; Zhang & Pfister, 2021; Shi et al., 2023).
@@
Cost-sensitive techniques. These techniques, including cost-sensitive learning and the incorporation of class weights assign different penalization costs to losses on different classes. They include cost-sensitive SVM (Iranmehr et al., 2019; Masnadi-Shirazi & Vasconcelos, 2010) and other cost-senstive methods (Elkan, 2001; Zhou & Liu, 2005; Zhao et al., 2018; Zhang et al., 2018, 2019; Sun et al., 2007; Fan et al., 2017; Jamal et al., 2020; Zhang et al., 2022; Wang et al., 2022; Xiao et al., 2023; Suh & Seo, 2023). The weights are often determined by the relative number of samples in each class or a notion of effective sample size (Cui et al., 2019).
@@
These two method categories are very related and can actually be shown to be equivalent in the limit. Cost-sensitive methods can be viewed as more efficient, flexible and principled techniques for implementing data sampling methods. However, these methods often risk overfitting the minority class or discarding valuable information from the majority class. Both methods inherently bias the input training data distribution and suffer from Bayes-inconsistency (in Section 6, we prove that cost-sensitive methods do not admit Bayes-consistency). While they have been both reported to be effective in various instances, this varies and depends on the problem, the distribution, the choice of predictors, and the performance metric adopted and they have been reported not to be effective in all cases (Van Hulse et al., 2007). Additionally, cost-sensitive methods often resort to careful tuning of hyperparameters. Hybrid approaches attempt to combine the strengths of data modification and cost-sensitive methods but often inherit their respective limitations.
@@
Logistic loss modifications. A family of more recent methods rely on logistic loss modifications. They consist of modifying the logistic loss by augmenting each logit (or predicted score) with an additive hyperparameter. They can be equivalently described as a cost-sensitive modification of the exponential terms appearing in the definition of the logistic loss. They include the Balanced Softmax loss (Jiawei et al., 2020), the Equalization loss (Tan et al., 2020), and the ldam loss (Cao et al., 2019). Other similar additive change methods use quadratically many hyperparameters with a distinct additive parameter for each pair of logits. They include the logit adjustment methods of Menon et al. (2021) and Khan et al. (2019). Menon et al. (2021) argue that their specific choice of the hyperparameter values is Bayes-consistent. A multiplicative modification of the logits, with one hyperparameter per class label is advocated by Ye et al. (2020). This can be equivalently viewed as normalizing scoring functions (or feature vectors in the linear case) beforehand, which is a standard method used in many learning applications, irrespective of the presence of imbalanced classes. The Vector-Scaling loss of Kini et al. (2021) combines the additive modification of the logits with this multiplicative change. These authors further present an analysis of this method in the case of linear predictors, underscoring the specific benefits of the multiplicative changes. As already pointed out, the multiplicative changes coincide with prior rescaling or renormalization of the feature vectors, however.
@@
Other methods. Additional approaches for tackling imbalanced datasets (see Zhang et al. (2023)) include post-hoc correction of decision thresholds (Fawcett & Provost, 1996; Collell et al., 2016) or weights (Kang et al., 2020; Kim & Kim, 2020)], as well as information and data augmentation via transfer learning, or distillation (Li et al., 2024b).
@@
First, most existing solutions are heuristic-driven and lack a solid theoretical foundation, making their performance difficult to predict across varying contexts.
In fact, we are not aware of any analysis of the generalization guarantees for these methods, with the exception of that of (Cao et al., 2019; Jiawei et al., 2020; Wang et al., 2023). However, as further discussed in Section 6, the analysis presented by these authors is limited to the balanced loss, which equalizes the impact of each class by weighting errors inversely to class frequency. More specifically, the analysis in (Cao et al., 2019) is limited to binary classification and holds only for the separable case. The balanced loss function differs from the target misclassification loss. It has been argued, and that is important, that the balanced loss admits beneficial fairness properties when class labels correlate with demographic attributes as it treats all class errors equally. The balanced loss is also the metric considered in the analysis of several of the logistic loss modifications papers (Cao et al., 2019; Menon et al., 2021; Ye et al., 2020; Kini et al., 2021). However, class labels do not alway relate to demographic attributes. Furthermore, many other criteria are considered for fairness purposes and in many machine learning applications, the misclassification remains the key target loss function to minimize. We will show that, even in the special case of the analysis of (Cao et al., 2019), the solution they propose is the opposite of the one corresponding to our theoretical analysis for the standard misclassification loss. We further show that their solution is empirically outperformed by ours.